OpenStudy (anonymous):

Find db/da for a^3b^2 = sin(a). Please help. :((

ganeshie8 (ganeshie8):

\(\large a^3b^2 = \sin(a)\)

ganeshie8 (ganeshie8):

\(a\) and \(b\) are variables

ganeshie8 (ganeshie8):

\(\large a^3b^2 = \sin(a)\) differentiate both sides with.respect.to \(a\)

OpenStudy (anonymous):

so would it be 3a^2b^2 = cos(a)?

ganeshie8 (ganeshie8):

you need to use product rule for left hand side

OpenStudy (anonymous):

got it

ganeshie8 (ganeshie8):

\(\large a^3b^2 = \sin(a) \) differentiate both sides with.respect.to \(a\) \(\large a^3(b^2)' + (a^3)'b^2 = \cos(a) \)

OpenStudy (anonymous):

(a^3)(2b)+(b^2)(3a^2) = cos (a) right?

ganeshie8 (ganeshie8):

not quite, b is a function of a here - b is dependent variable, so by chain rule, u need to differentiate it again

ganeshie8 (ganeshie8):

\(\large a^3(b^2)' + (a^3)'b^2 = \cos(a)\) \(\large a^3(2b) b' + (3a^2)b^2 = \cos(a)\)

ganeshie8 (ganeshie8):

solve \(\large b'\)

ganeshie8 (ganeshie8):

btw, \(\large b'\) is same as \(\large \frac{db}{da}\)

OpenStudy (anonymous):

so would it be -a^3(2b)/(3a^2)b^2 +cos(a)

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

I prefer to just gave up haha. I can't think anymore but okay..

ganeshie8 (ganeshie8):

\(\large a^3(2b) b' + (3a^2)b^2 = \cos(a)\) \(\large a^3(2b) b' = \cos(a) - (3a^2)b^2\)

ganeshie8 (ganeshie8):

dont give up lol^ we're done wid calc :) its just algebra, solve \(\large b'\)

OpenStudy (anonymous):

i think i got it. waitt

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):

-cos(a)-3a^2b^2/ (a^32b)?

ganeshie8 (ganeshie8):

correct ! except for that negative sign..

ganeshie8 (ganeshie8):

\(\large a^3(2b) b' + (3a^2)b^2 = \cos(a)\) \(\large a^3(2b) b' = \cos(a) - (3a^2)b^2\) \(\large a^3(2b) b' + (3a^2)b^2 = \cos(a)\) \(\large b' = \frac{\cos(a) - (3a^2)b^2}{a^3(2b)}\)

ganeshie8 (ganeshie8):

or may be put it like this : \(\large b' = \frac{\cos(a) - 3a^2b^2}{2a^3b}\)

OpenStudy (anonymous):

oh haha that was supposed to be an equal sign.

OpenStudy (anonymous):

but thank you :DDD.

ganeshie8 (ganeshie8):

cool :) u wlc !